sadie5:

Which equation is true for the value x = 15? A. 2(x + 3) = 40 B. 2(x − 5) = 30 C. 2(x + 5) = 40 D. x + 2x = 30 E. 3x − x = 45

sadie5:

what do u think

KynxKatz:

What do you know?

supie:

Use the method that was shown on one of your previous posts: plug in 15 for x in the answer choices then solve it and see if it's true \(\bf A.~2(15 + 3) = 40 \\B.~2(15 − 5) = 30\\ C.~2(15 + 5) = 40\\ D.~15 + 2(15) = 30\\ E.~3(15) − 15 = 45 \)

QuestionCoveBot:

And then simplify each equation.

sadie5:

what is the correct answer

sadie5:

i think it C

supie:

I think you're right

sadie5:

🌺

sadie5:

thanks